OpenStudy (anonymous):

What is the geometric mean between 2 and 50?

OpenStudy (anonymous):

to find the geometric mean, you multiply the numbers and take the square root: \[\sqrt{2*50} = \sqrt{100} = \sqrt{10^2} = 10\]

OpenStudy (amistre64):

well, not the square root; but the n-root where n is the number of terms you have. In this case its 2 terms so square root it :)

OpenStudy (anonymous):

ah, i stand corrected. thank you :)